OpenStudy (anonymous):

Would 2x-3y=8 -5x+2y=13 give me -5 for the x coordinate?

OpenStudy (anonymous):

\[2x-3y=8[1]\] \[-5x+2y=13[2]\] \[[1]\times 2\] \[2(2x-3y)=2(8)\] \[4x-6y=16[3]\] \[[2]\times 3\] \[3(-5x+2y)=3(13)\] \[-15x+6y=39[4]\] \[[3]+[4]\] \[(4x-6y)+(-15x+6y)=(16)+(39)\] \[4x-15x=55\] \[-11x=55\] \[x=-5\] So yes, you're correct.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No worries.